# The OpenHands Software Agent SDK: A Composable and Extensible Foundation for Production Agents

### v1

@@ Abstract
Agents are now used widely in the process of software development, but building production-ready software engineering agents is a complex task. Deploying software agents effectively requires flexibility in implementation and experimentation, reliable and secure execution, and interfaces for users to interact with agents. In this paper, we present the OpenHands Software Agent SDK, a toolkit for implementing software development agents that satisfy these desiderata. This toolkit is a complete architectural redesign of the agent components of the popular OpenHands framework for software development agents, which has 64k+ GitHub stars.
To achieve flexibility, we design a simple interface for implementing agents that requires only a few lines of code in the default case, but is easily extensible to more complex full-featured agents with features such as custom tools, memory management, and more. For security and reliability, it delivers seamless local-to-remote execution portability, integrated REST/WebSocket services. For interaction with human users, it can connect directly to a variety of interfaces, such as visual workspaces (VS Code, VNC, browser), command-line interfaces, and APIs. Compared with existing SDKs from OpenAI, Claude and Google, OpenHands uniquely integrates native sandboxed execution, lifecycle control, model-agnostic multi-LLM routing, and built-in security analysis. Empirical results on SWE-Bench Verified and GAIA benchmarks demonstrate strong performance. Put together, these elements allow the OpenHands Software Agent SDK to provide a practical foundation for prototyping, unlocking new classes of custom applications, and reliably deploying agents at scale.
@@ 1 Introduction
In software engineering, AI agents have evolved from assistive tools (GitHub, 2021; Cursor Team, 2024) to autonomous systems capable of hours-long async execution on complex tasks (e.g., Devin (Cognition AI, 2024), Claude Code (Anthropic, 2025b), OpenHands (Wang et al., 2025)). Reliably deploying autonomous agents in production requires system foundations—including durable state management, safe execution in a sandbox, and consistent behavior across environments ranging from local to containerized cloud deployments—that were unnecessary for earlier assistive tools that rely on largely local user-driven workflows.
OpenHands (Wang et al., 2025) demonstrated that open-source software agents can achieve broad adoption and contributions, reaching over 64k GitHub stars and hundreds of contributors in just 18 months (Neubig, 2025). As the project scaled, the original monolithic architecture – what we refer to as OpenHands V0 – exposed growing architectural tensions. The early design was driven by the need for fast prototyping and iteration, combining agent logic, evaluation, and applications in a single codebase. Over time, this approach led to rigid sandboxing assumptions, sprawling mutable configuration, and tight coupling between research and production, eventually making a full architectural redesign unavoidable.
@@
Building on these principles, we introduce OpenHands V1, a complete software-agent ecosystem—including CLI and GUI applications—built on a shared foundation: the OpenHands Software Agent SDK (Fig. 1(b)). The SDK defines an event-sourced state model with deterministic replay, an immutable configuration for agents, and a typed tool system with MCP integration. Its workspace abstraction enables the same agent to run locally for prototyping or remotely in secure, containerized environments with minimal code changes. Unlike prior library-only SDKs (Anthropic, 2025a; OpenAI, 2024), OpenHands includes a built-in REST/WebSocket server for remote execution and a suite of interactive workspace interfaces—a browser-based VSCode IDE, VNC desktop, and persistent Chromium browser—for human inspection and control.
We systematically compare 31 features of our SDK with those of the OpenAI Agents SDK, Claude Agent SDK, and Google ADK, and find that, although 15 features are shared with at least one of them, our SDK uniquely combines 16 additional features, including native remote execution, a production server with sandboxing, and model-agnostic multi-LLM routing across 100+ providers. Our SDK further adds a security analyzer for agent actions, flexible lifecycle control (pause/resume, sub-agent delegation, history restore, etc), and built-in QA instrumentation (unit tests, LLM-based integration tests, and evaluation benchmarks) for production reliability.
Across multiple LLM backends, our SDK achieves strong results on SWE-Bench Verified and GAIA benchmarks, demonstrating both state-of-the-art performance and the generality of the architecture to operate consistently across diverse model providers.
The OpenHands Software Agent SDK is fully open-sourced under the MIT License at https://github.com/OpenHands/software-agent-sdk.
@@ 2.1 Agent Design
Modern AI agents can be viewed as systems that perceive, reason, and act within an environment to accomplish goals over time (Russell and Norvig, 2009). In recent work (Wang et al., 2023), agent design typically centers on these interconnected components:
@@ 2.2 Software Agents
In software engineering contexts, these agentic components manifest concretely in systems like OpenHands (Wang et al., 2025), which emulate a developer’s end-to-end workflow. The environment is a sandboxed workspace exposing a filesystem, terminal, and web interface, allowing the agent to observe program output and test feedback. Its state and memory are represented through an event log that records commands, edits, and results, providing persistent context across actions. The agent acts by editing files, running tests, or invoking structured tools like web browser to interact with its environment.
@@ 3 Challenges and Design Principles
As OpenHands evolved, it has accumulated significant architectural complexity: core agent logic, multiple CLIs, a web server for the GUI, frontend code, and diverse runtime providers (e.g., local Docker, Kubernetes in production) all coexist in a single codebase. Much of this architecture dated back to the project’s first weeks, before the introduction of LLM structured tool use or the Model Context Protocol (MCP; MCP Team 2025). This section revisits the key architectural tensions observed in OpenHands V0 and distills them into four design principles guiding V1.
@@ 4.2 Event-Sourced State Management
Event Hierarchy. The event system uses a multi-level hierarchy shown in Table 1. At the base, Event provides immutable structure (ID, timestamp, source) with type-safe serialization via discriminated unions (Pydantic Team, 2025). LLMConvertibleEvent adds to_llm_message() for converting events into LLM format. The action-observation loop uses ActionEvent for tool calls and ObservationBaseEvent subclasses for results – all these are subclasses of LLMConvertibleEvent. Internal events (condensation, state updates, etc) inherit directly from Event for bookkeeping without LLM exposure.
@@ 4.6 Context Window Management
Condensers enable long-running conversations and, as a mechanism for constraining the tokens consumed per step, can reduce the overall cost of a conversation. LLMSummarizingCondenser (the default condenser) has been shown to reduce API costs by up to $2\times$ with no degradation in agent performance (Smith, 2025).
@@ 5 Reliability and Evaluation
We assess the reliability and performance of the OpenHands Agent SDK through two complementary processes: continuous testing and benchmark evaluation. The continuous testing pipeline combines both programmatic and LLM-based tests and runs automatically on every pull request and once per day. It checks that the SDK behaves consistently across multiple language models, catching regressions in reasoning, tool use, and state management early. These automated tests cost only \$0.5–\$3 per full run and complete in less than 5 minutes. In parallel, benchmark evaluations measure the SDK’s overall capabilities on standardized agentic tasks, providing a broader view of model quality and system performance.
@@ 5.1 Continuous Quality Assurance
5.2 Benchmarks
The SDK offers built-in support for a diverse range of academic benchmarks evaluating agentic capability.
Benchmark Model Performance
SWE-Bench (Jimenez et al., 2024) Claude Sonnet 4.5 $72.8\%$
Claude Sonnet 4 $68.0\%$
GPT-5 (reasoning=high) $68.8\%$
Qwen3 Coder 480B A35B $65.2\%$
GAIA (Mialon et al., 2023a) Claude Sonnet 4 $57.6\%$
Claude Sonnet 4.5 $67.9\%$
GPT-5 (reasoning=high) $62.4\%$
Qwen3 Coder 480B A35B $41.2\%$
Table 2: Benchmark performance of OpenHands Software Agent SDK. Evaluations were performed at commit 54c5858 of the SDK and commit 88f1d80 of the benchmarks.
As shown in Tab. 2, the SDK demonstrates competitive performance on software engineering and general agent benchmarks.
On SWE-Bench Verified (Jimenez et al., 2024) that measures an agent’s ability in software engineering tasks, SDK achieves 72% resolution rate using Claude Sonnet 4.5 with extended thinking; On the GAIA (val set, Mialon et al. (2023b)) that measuresan agent’s generic computer task-solving capability, SDK achieves 67.9% accuracy with Claude Sonnet 4.5, demonstrating effective multi-step reasoning and tool use. Additionally, Qwen3 Coder 480B (Yang et al., 2025), a strong open model for coding, also stays at a competitive score of 41.21%. These results are slightly better than those of OpenHands-Versa (Soni et al., 2025), validating that the SDK’s architecture does not sacrifice agentic capability and maintains performance competitive with research-focused systems.
OpenAI Agents SDK Claude Agent SDK Google Agent Development Kit OpenHands Software Agent SDK
@@ 5.2 Benchmarks
MCP Support $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Custom Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
History Persistence & Restore $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Sub-agent Delegation $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Model Agnostic (100+ LLMs) $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Multi-LLM Routing $\times$ $\times$ $\times$ $\checkmark$
Conversation Cost & Token Tracking $\checkmark$ $\checkmark$ $\times$ $\checkmark$
Pause/Resume Agent Execution $\checkmark$ $\times$ $\times$ $\checkmark$
Native Support for Non-function-calling models $\times$ $\times$ $\times$ $\checkmark$
Security Analyzer for Agent Action $\times$ $\times$ $\checkmark$ $\checkmark$
Action Confirmation Policies $\sim$ $\times$ $\sim$ $\checkmark$
Context Files (e.g., repo.md, AGENTS.md) $\times$ $\checkmark$ $\times$ $\checkmark$
Agent Skills $\times$ $\checkmark$ $\times$ $\checkmark$
Context Condensation $\times$ $\checkmark$ $\times$ $\checkmark$
TODO List Planner $\times$ $\checkmark$ $\times$ $\checkmark$
Tmux-based Interactive Bash Terminal $\times$ $\times$ $\times$ $\checkmark$
Auto-Generated Conversation Titles $\times$ $\times$ $\times$ $\checkmark$
Secrets Management with Auto-Masking $\times$ $\times$ $\times$ $\checkmark$
Agent Stuck Detection $\times$ $\times$ $\times$ $\checkmark$
Long-term Memory across Sessions $\times$ $\times$ $\checkmark$ $\times$
Production Server Features
Builtin REST+WebSocket Server $\times$ $\times$ $\times$ $\checkmark$
Session-Based Authentication $\times$ $\times$ $\times$ $\checkmark$
Builtin Remote Agent Execution $\times$ $\times$ $\times$ $\checkmark$
Agent Environment Sandboxing $\times$ $\times$ $\sim$ $\checkmark$
VNC Desktop for Agent Workspace $\times$ $\times$ $\times$ $\checkmark$
VSCode Web for Agent Workspace $\times$ $\times$ $\times$ $\checkmark$
Builtin Chromium Browser for Agent $\times$ $\times$ $\times$ $\checkmark$
Continous Quality Assurance
Programmatic tests $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Strong Type Checking $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Frequent LLM-based tests $\times$ $\checkmark$ $\checkmark$ $\checkmark$
Built-in Academic Benchmark Evaluation $\times$ $\times$ $\times$ $\checkmark$
Table 3: Agent SDK features comparisons ($\checkmark$= full support, $\sim$= partial, $\times$= absent; as of Oct 29, 2025). OpenHands software agent SDK uniquely combines: (i) native remote execution with secure sandboxing (vs. external orchestration for OpenAI and local-only for Claude), (ii) a built-in production server with REST+WebSocket APIs and workspace access (vs. library-only SDKs), (iii) LLM-powered action-level security analysis (vs. manual or external guardrails), and (iv) model-agnostic multi-LLM routing with first-class support for non-function-calling models.

### v2

@@ Abstract
Agents are now used widely in the process of software development, but building production-ready software engineering agents is a complex task. Deploying software agents effectively requires flexibility in implementation and experimentation, reliable and secure execution, and interfaces for users to interact with agents. In this paper, we present the OpenHands Software Agent SDK, a toolkit for implementing software development agents that satisfy these desiderata. This toolkit is a complete architectural redesign of the agent components of the popular OpenHands framework for software development agents.
To achieve flexibility, we design a simple interface for implementing agents that requires only a few lines of code in the default case, but is easily extensible to more complex full-featured agents with features such as custom tools, memory management, and more. For security and reliability, it delivers seamless local-to-remote execution portability, integrated REST/WebSocket services. For interaction with human users, it can connect directly to a variety of interfaces, such as visual workspaces (VS Code, VNC, browser), command-line interfaces, and APIs. Compared with existing SDKs from OpenAI, Claude and Google, OpenHands uniquely integrates native sandboxed execution, lifecycle control, model-agnostic multi-LLM routing, and built-in security analysis. We validate the architecture empirically: production deployment data shows that V1 substantially reduces system-attributable failures over V0 with negligible event-sourcing overhead, and evaluations across multiple models and benchmarks demonstrate strong agent performance. Put together, these elements allow the OpenHands Software Agent SDK to provide a practical foundation for prototyping, unlocking new classes of custom applications, and reliably deploying agents at scale.
[…]
@@ 1 Introduction
In software engineering, AI agents have evolved from assistive tools (copilot2021github; cursor2024) to autonomous systems capable of hours-long async execution on complex tasks (e.g., Devin (devin2024cognition), Claude Code (claudecode), OpenHands (wang2025openhands)). Reliably deploying autonomous agents in production requires system foundations—including durable state management, safe execution in a sandbox, and consistent behavior across environments ranging from local to containerized cloud deployments—that were unnecessary for earlier assistive tools that rely on largely local user-driven workflows. Yet no consensus exists on how to architect these foundations: existing SDKs from major providers differ widely in isolation models, state management, and deployment assumptions, leaving practitioners without a well-validated reference design.
OpenHands (wang2025openhands) demonstrated that open-source software agents can achieve broad adoption and contributions, reaching over 64k GitHub stars and hundreds of contributors in just 18 months (neubig2025-one-year-of). As the project scaled, the original monolithic architecture—what we refer to as OpenHands V0—exposed growing architectural tensions. The early design was driven by the need for fast prototyping and iteration, combining agent logic, evaluation, and applications in a single codebase. Over time, this approach led to rigid sandboxing assumptions, sprawling mutable configuration, and tight coupling between research and production, eventually making a full architectural redesign unavoidable.
@@
Building on these principles, we introduce OpenHands V1, a complete software-agent ecosystem—including CLI and GUI applications—built on a shared foundation: the OpenHands Software Agent SDK (Fig. 1(b)). We present the SDK as a reference architecture for production agent systems, codifying lessons from 18 months of open-source development and production deployment into an empirically validated design. The SDK defines an event-sourced state model with deterministic replay, an immutable configuration for agents, and a typed tool system with MCP integration. Its workspace abstraction enables the same agent to run locally for prototyping or remotely in secure, containerized environments with minimal code changes. Unlike prior library-only SDKs (anthropic-agents; openai_agents_2024), OpenHands includes a built-in REST/WebSocket server for remote execution and a suite of interactive workspace interfaces for human inspection and control.
@@
We systematically compare features of our SDK with those of the OpenAI Agents SDK, Claude Agent SDK, Google ADK, and LangChain (Tab. 6), finding that our SDK uniquely combines native remote execution with sandboxing, a production server, model-agnostic multi-LLM routing across 100+ providers, a security analyzer for agent actions, and built-in QA instrumentation for production reliability.
We support this architectural contribution with empirical evidence at two levels. At the systems level, a 15-day production comparison shows V1 reduces system-attributable failures by 61% relative to V0, and event-sourcing overhead is negligible. At the agent capability level, evaluations across 14 language models and five benchmark categories—including SWE-Bench Verified, GAIA, SWE-Bench Multimodal, SWT-Bench, and Commit0—demonstrate that the redesigned architecture maintains strong agent performance.
The OpenHands Software Agent SDK [^1] https://github.com/OpenHands/software-agent-sdk and evaluation harness [^2] https://github.com/OpenHands/benchmarks are fully open-sourced under the MIT License.
@@ 2.1 Agent Design
Modern AI agents can be viewed as systems that perceive, reason, and act within an environment to accomplish goals over time (10.5555/1671238). In recent work (Wang2023ASO), agent design typically centers on these interconnected components:
@@ 2.2 Software Agents
In software engineering contexts, these agentic components manifest concretely in systems like OpenHands (wang2025openhands), which emulate a developer’s end-to-end workflow. The environment is a sandboxed workspace exposing a filesystem, terminal, and web interface, allowing the agent to observe program output and test feedback. Its state and memory are represented through an event log that records commands, edits, and results, providing persistent context across actions. The agent acts by editing files, running tests, or invoking structured tools like web browser to interact with its environment.
@@ 3 Challenges and Design Principles
As OpenHands evolved, it has accumulated significant architectural complexity: core agent logic, multiple CLIs, a web server for the GUI, frontend code, and diverse runtime providers (e.g., local Docker, Kubernetes in production) all coexist in a single codebase. Much of this architecture dated back to the project’s first weeks, before the introduction of LLM structured tool use or the Model Context Protocol (MCP; mcp2025intro). This section revisits the key architectural tensions observed in OpenHands V0 and distills them into four design principles guiding V1.
@@ 4.2 Event-Sourced State Management
Event Hierarchy. The event system uses a multi-level hierarchy shown in Table 1. At the base, Event provides immutable structure (ID, timestamp, source) with type-safe serialization via discriminated unions (pydantic_discriminated_unions). LLMConvertibleEvent adds to_llm_message() for converting events into LLM format. The action-observation loop uses ActionEvent for tool calls and ObservationBaseEvent subclasses for results – all these are subclasses of LLMConvertibleEvent. Internal events (condensation, state updates, etc) inherit directly from Event for bookkeeping without LLM exposure.
@@
Empirical measurements on agent traces confirm that event-sourcing overhead is negligible relative to LLM round-trip times, with sub-millisecond persist latency and crash recovery under 20 ms (see §5.2 for details).
@@ 4.6 Context Window Management
Condensers enable long-running conversations and, as a mechanism for constraining the tokens consumed per step, can reduce the overall cost of a conversation. LLMSummarizingCondenser (the default condenser) has been shown to reduce API costs by up to $2\times$ with no degradation in agent performance (smith2025-openhands-context-condensensation).
@@ 5 Reliability and Evaluation
We assess the OpenHands Agent SDK through four complementary lenses: (1) production reliability comparing V0 and V1, (2) systems overhead of the event-sourcing mechanism, (3) continuous testing, and (4) benchmark evaluation across diverse models and tasks.
5.1 Production Reliability: V0 vs. V1
To directly evaluate whether the architectural redesign improves operational reliability, we analyzed production logs from a 15-day parallel rollout in which V0 and V1 served users simultaneously. Exceptions were extracted from conversation failure logs and categorized into three classes (Tab. 7):
To ensure comparability, we report system-attributable errors—error classes directly attributable to agent infrastructure or SDK logic, excluding LLM provider errors that are external to both architectures. Results are shown in Tab. 2.
Table 2: System-attributable errors per 1k conversations during 15-day production rollout.
V0 (Legacy) V1 (SDK)
Infrastructure errors 69.8 / 1k 0.0 / 1k
SDK errors N/A 29.7 / 1k
System-attributable error rate 78.0 / 1k 30.0 / 1k
Result: V1 reduces system-attributable failures by 61% (from 78.0 to 30.0 errors per 1k conversations). The eliminated V0 infrastructure errors were dominated by: HTTPStatusError (401) at 43.0 / 1k (inter-pod authentication failures), AgentRuntimeNotReadyError at 18.8 / 1k (runtime pod readiness races), and connection/timeout errors at 3.1 / 1k (network instability). These failures stem from V0’s inter-pod HTTP communication between the conversation manager and execution runtime. V1’s co-located execution model eliminates these observed failure modes by removing this dependency entirely. The remaining V1 SDK errors (29.7 / 1k) were dominated by a condensation bug discovered during the rollout of extended thinking support, as the LLM provider introduced additional constraints on the event/message interface. This issue has since been fixed in the official release.
5.2 Event-Sourcing Systems Overhead
To quantify the I/O cost of event sourcing, we replayed real payloads from 433 SWE-Bench Verified conversations (39,870 events) through the production LocalFileStore path (Tab. 3). All persist and recovery latencies are negligible relative to LLM round-trip times (typically 1–30 s), storage grows linearly, and crash recovery completes in under 20 ms even for the longest observed conversation (358 events). These results demonstrate that event sourcing adds minimal overhead to the agent execution loop.
Table 3: Event-sourcing systems overhead measured on SWE-Bench Verified traces (433 conversations, 39,870 events).
Metric Median P95 At max (358 events)
Per-event persist latency 0.20 ms 0.31 ms
Action cycle persist (Action+Obs) 0.40 ms 0.56 ms
Full state replay 4.1 ms 9.7 ms 18.9 ms
Crash recovery (replay + unmatched-action scan) 7.4 ms 14.9 ms 32.1 ms
Storage per conversation 380 KB 1.4 MB 3.4 MB
@@ 5.3 Continuous Quality Assurance
5.4 Benchmark Evaluation
We validate the SDK on academic benchmarks along two axes: first, that the V0-to-V1 architectural transition preserves (and, where the architecture enables new provider features, improves) agentic capability; second, that the SDK generalizes across models and task types.
The evaluation spans five categories of software engineering work: Issue Resolution (SWE-Bench Verified (jimenez2024swebench)), Greenfield Development (Commit0 (commit0)), Frontend Development (SWE-Bench Multimodal (yang2024swebenchmultimodal)), Software Testing (SWT-Bench (mundler2024swtbench)), and Information Gathering (GAIA (mialon2023gaiabenchmarkgeneralai)). Results for all 14 evaluated models are maintained on the continuously updated OpenHands Index. [^3] https://index.openhands.dev
V0 vs. V1: Capability Preservation. We compare V0 and V1 on SWE-Bench Verified using matched models to isolate the contribution of the SDK redesign (Tab. 4). With Claude Sonnet 4, V0 and V1 achieve identical performance ($68.0\%$), confirming that the architectural redesign preserves baseline agentic capability. With Sonnet 4.5, V1 gains +8.2+8.2 points; we attribute this to extended thinking support, which V1’s event-sourced architecture integrates naturally but would require significant engineering effort to retrofit into V0’s multi-component design.
Model V0 V1 (SDK)
Claude Sonnet 4.5 $64.6\%$ $\mathbf{72.8\%}$
Claude Sonnet 4 $\% 68.068.0\%$
Table 4: V0 vs. V1 on SWE-Bench Verified with matched models. Sonnet 4 confirms capability parity; the Sonnet 4.5 gain is attributed to extended thinking support enabled by V1’s architecture.
Comprehensive Multi-Model Evaluation. We evaluated 14 language models—7 closed-source (Anthropic, OpenAI, Google) and 7 open-weights (MiniMax, DeepSeek, Zhipu AI, Moonshot AI, Alibaba, NVIDIA)—across all five categories. Tab. 5 summarizes the best SDK result per category alongside published state-of-the-art. The SDK achieves SOTA on 3 of 5 benchmarks and remains competitive on the other two (within $2.6$ and $5.2$ points), using a single model per evaluation while several SOTA systems employ multi-model orchestration.
Issue Resol. Greenfield Frontend Testing Info. Gather.
SWE-Bench V. Commit0 SWE-Bench MM SWT-Bench V. GAIA (test)
Published SOTA $79.2\%$ $12.5\%$ - [^4] OpenHands Index uses the test dataset of SWE-bench Multimodal, for which no state-of-the-art results have yet been published. $84.0\%$ $74.6\%$
Best SDK $76.6\%$ $\mathbf{56.2\%}$ $\mathbf{44.1\%}$ $78.8\%$ $\mathbf{80.0\%}$
Model Opus 4.5 GPT-5.4 Gemini 3.1 Pro Opus 4.6 Opus 4.6
2nd SDK $75.6\%$ $56.2\%$ $41.8\%$ $5\% 78.87$$8\%$
Model GPT-5.4 Opus 4.6 Opus 4.6 Opus 4.5 GPT-5.4
Table 5: Multi-model benchmark evaluation across five SE task categories (14 models evaluated). Bold indicates the SDK exceeds published SOTA. Full per-model results are available on the OpenHands Index (https://index.openhands.dev).
SDK-Enabled Insights. The uniform evaluation harness enables observations that would be difficult without a model-agnostic SDK. Models exhibit clear task-specific specialization: Claude models dominate issue resolution and testing, while GPT-5.4 leads in long-horizon greenfield development ($62.5\%$, +12.5+12.5 pts ahead of second place).
OpenAI Agents SDK Claude Agent SDK Google Agent Dev. Kit LangChain/ LangGraph OpenHands Agent SDK
@@ 5.4 Benchmark Evaluation
MCP Support $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Custom Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
History Persistence & Restore $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Sub-agent Delegation $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Model Agnostic (100+ LLMs) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
Multi-LLM Routing $\times$ $\times$ $\times$ $\sim$ $\checkmark$
Conversation Cost & Token Tracking $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Pause/Resume Agent Execution $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$
Native Support for Non-function-calling models $\times$ $\times$ $\times$ $\times$ $\checkmark$
Security Analyzer for Agent Action $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$
Action Confirmation Policies $\sim$ $\times$ $\sim$ $\sim$ $\checkmark$
Context Files (e.g., repo.md, AGENTS.md) $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$
Agent Skills $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$
Context Condensation $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Secrets Management with Auto-Masking $\times$ $\times$ $\times$ $\times$ $\checkmark$
Agent Stuck Detection $\times$ $\times$ $\times$ $\times$ $\checkmark$
Production Server & Deployment
Builtin REST+WebSocket Server $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
Session-Based Authentication $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
Builtin Remote Agent Execution $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
Agent Environment Sandboxing $\times$ $\times$ $\sim$ $\times$ $\checkmark$
Continuous Quality Assurance
Programmatic tests $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Strong Type Checking $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Frequent LLM-based tests $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$
Built-in Academic Benchmark Evaluation $\times$ $\times$ $\times$ $\times$ $\checkmark$
Table 6: Feature comparison across major agent SDKs ($\checkmark$= full support, $\sim$= partial, $\times$= absent). Methodology: Features were assessed from official documentation, public repositories, and release notes as of October 2025. Versions assessed: OpenAI Agents SDK v0.4.2, Claude Agent SDK v0.1.6, Google ADK v1.17.0, LangChain v1.0.3 / LangGraph v1.0.2, OpenHands Agent SDK v1.0.0. “Partial” indicates the capability requires significant external setup or community extensions rather than built-in support. OpenHands uniquely combines: (i) native remote execution with environment sandboxing, (ii) LLM-powered action-level security analysis, (iii) model-agnostic multi-LLM routing with first-class support for non-function-calling models, and (iv) built-in academic benchmark evaluation for continuous quality assurance.